# From Hypothesis to Publication: A Comprehensive Survey of AI-Driven Research Support Systems

### v1

@@ Abstract
Research is a fundamental process driving the advancement of human civilization, yet it demands substantial time and effort from researchers. In recent years, the rapid development of artificial intelligence (AI) technologies has inspired researchers to explore how AI can accelerate and enhance research. To monitor relevant advancements, this paper presents a systematic review of the progress in this domain. Specifically, we organize the relevant studies into three main categories: hypothesis formulation, hypothesis validation, and manuscript publication. Hypothesis formulation involves knowledge synthesis and hypothesis generation. Hypothesis validation includes the verification of scientific claims, theorem proving, and experiment validation. Manuscript publication encompasses manuscript writing and the peer review process. Furthermore, we identify and discuss the current challenges faced in these areas, as well as potential future directions for research. Finally, we also offer a comprehensive overview of existing benchmarks and tools across various domains that support the integration of AI into the research process. We hope this paper serves as an introduction for beginners and fosters future research. Resources have been made publicly available at https://github.com/zkzhou126/AI-for-Research.
@@
Zekun Zhou1, Xiaocheng Feng1,2thanks: Corresponding Author, Lei Huang1, Xiachong Feng3, Ziyun Song1, Ruihan Chen1, Liang Zhao1, Weitao Ma1, Yuxuan Gu1, Baoxin Wang4, Dayong Wu4, Guoping Hu4, Ting Liu1,2, Bing Qin1,2 1 Harbin Institute of Technology, Harbin, China  2 Peng Cheng Laboratory, Shenzhen, China 3 The University of Hong Kong, China  4 iFLYTEK Research, China {zkzhou,xcfeng,lhuang,zysong,rhchen,lzhao,wtma,yxgu,tliu,qinb}@ir.hit.edu.cn fengxc@hku.hk{bxwang2,dywu2,gphu}@iflytek.com
@@ 1 Introduction
Comparing with Existing Surveys Although Luo et al. (2025) reviews the application of AI in research, it predominantly focuses on LLMs, while neglecting the knowledge synthesis that precedes hypothesis generation and the theoretical validation of hypothesis. Other surveys concentrate on more specific areas, such as paper recommendation (Beel et al., 2016; Bai et al., 2019; Kreutz and Schenkel, 2022), scientific literature review (Altmami and Menai, 2022), scientific claim verification (Vladika and Matthes, 2023; Dmonte et al., 2024), theorem proving (Li et al., 2024e), manuscript writing (Li and Ouyang, 2024), and peer review  (Lin et al., 2023a; Kousha and Thelwall, 2024). Additionally, certain surveys emphasize the application of AI in scientific domains (Zheng et al., 2023; Zhang et al., 2024d).
Contributions Our contributions can be summarized as follows: (1) We align the relevant fields with the research process of human researchers, systematically integrating and extending these aspects while primarily focusing on the research process itself. (2) We introduce a meticulous taxonomy (shown in Figure 2). (3) We provide a summary of tools that can assist in the research process. (4) We formally define AI for research and clearly distinguish it from AI for science in  §A.
@@
Formulation  (§2) [ Knowledge Synthesize  (§2.1) [ Research Paper
@@
Survey Organization We first elaborate hypothesis formulation (§2), followed by hypothesis validation (§3) and manuscript publication (§4). Additionally, we present benchmarks (§5), and tools (§6) that utilized in research. Finally, we outline challenges as well as future directions (§7) and give a further discussion about open questions (§A).
@@ 2.1.2 Systematic Literature Review
Systematic literature review (SLR) constitutes a rigorous and structured methodology for evaluating and integrating prior research on a specific topic  (Webster and Watson, 2002; Zhu et al., 2023; Bolaños et al., 2024). In contrast to single-document summaries, SLR entails synthesizing information across multiple related scientific documents (Altmami and Menai, 2022). SLR can further be divided into two stages: outline generation and full-text generation (Shao et al., 2024; Agarwal et al., 2024b; Block and Kuckertz, 2024).
@@ 2.2 Hypothesis Generation
Hypothesis quality improvement has been addressed through feedback and iteration, as proposed by HypoGeniC  (Zhou et al., 2024) and MOOSE  (Yang et al., 2024b). Feedback mechanisms include direct responses to hypothesis  (Baek et al., 2024), experimental outcome evaluations  (Ma et al., 2024; Yuan et al., 2025), and automated peer review comments  (Lu et al., 2024). Beyond iterative feedback, collaborative efforts among researchers have also been recognized, leading to multi-agent hypothesis generation approaches  (Nigam et al., 2024; Ghafarollahi and Buehler, 2024). VIRSCI  (Su et al., 2024) further optimized this process by customizing knowledge for each agent. Additionally, the Nova framework  (Hu et al., 2024a) was introduced to refine hypothesis by leveraging outputs from other research as input.
@@ 3 Hypothesis Validation
In scientific research, any proposed hypothesis must undergo rigorous validation to establish its validity. As illustrated in Figure  4, this section explores the application of AI in verifying scientific hypothesis through three methodological components: Scientific Claim Verification, Theorem Proving, and Experiment Validation.
@@ 3.1 Scientific Claim Verification
In the verification results generation phase, studies propose leveraging LLMs to synthesize evidence into comprehensive information  (Kao and Yen, 2024; Cao et al., 2024b). FactKG  (Kim et al., 2023) and  Muharram and Purwarianti (2024) structured evidence into knowledge graphs, enabling claim attribution  (Dammu et al., 2024; Wu et al., 2023). Furthermore,  Atanasova et al. (2020); Krishna et al. (2022); Pan et al. (2023a); Eldifrawi et al. (2024); Zhang et al. (2024b) advocated for generating explanatory annotations alongside experimental outcomes during the verification process. Meanwhile,  Das et al. (2023); Altuncu et al. (2023) emphasized the critical role of domain expertise in ensuring accurate verification.
@@ 3.2 Theorem Proving
Theorem proving constitutes a subtask of logical reasoning, aimed at reinforcing the validity of the underlying theory within a hypothesis  (Yang et al., 2023c; Li et al., 2024e).
@@ 4.1 Manuscript Writing
A subset of research on AI in scientific writing has focused on citation text generation, which addresses the academic need for referencing prior work while mitigating model inaccuracies  (Gao et al., 2023b; Gu and Hahnloser, 2023). For instance,  Wang et al. (2022b) developed an automated citation generation system by integrating manuscript content with citation graphs. However, its reliance on rigid template-based architectures led to inflexible citation formats. This limitation motivated subsequent studies to propose incorporating citation intent as a control parameter during text generation, aiming to improve contextual relevance and rhetorical adaptability  (Yu et al., 2022; Jung et al., 2022; Gu and Hahnloser, 2024).
@@
In contrast to citation text generation, several studies have focused on related work generation in scholarly writing, emphasizing the production of multiple citation texts and the systematic analysis of inter-citation relationships  (Li and Ouyang, 2022, 2024). The ScholaCite framework  (Martin-Boyle et al., 2024) leveraged GPT-4 to cluster citation sources and generate draft literature review sections, although it required manually provided reference lists. By contrast, the UR3WG system  (Shi et al., 2023) adopted a retrieval-augmented architecture integrated with large language models to autonomously acquire relevant references. To improve the quality of generated related work sections,  Yu et al. (2024b) utilized graph neural networks (GNNs) to model complex relational dynamics between target manuscripts and cited literature, while  Nishimura et al. (2024) initiative advocated for explicit novelty assertions regarding referenced publications.
@@
The aforementioned investigations primarily focused on specific components of scientific writing, while a study by  Lai et al. (2024) explored the progressive generation of complete manuscripts via structured workflows. The AI-Scientist system  (Lu et al., 2024) further introduced section-wise self-reflection mechanisms to enhance compositional coherence. Several studies emphasized human-AI collaborative frameworks for improving writing efficiency  (Lin, 2024; Feng et al., 2024; Ifargan et al., 2024), whereas  Tang et al. (2024a) concentrated on enabling personalized content generation in multi-author collaborative environments. Following initial manuscript drafting, subsequent text revision processes were systematically examined  (Jourdan et al., 2023). The OREO system  (Li et al., 2022) utilized attribute classification for iterative in-situ editing, while  Du et al. (2022); Pividori and Greene (2024) incorporated researcher feedback loops for progressive text optimization. Notably,  Chamoun et al. (2024); D’Arcy et al. (2024b) proposed replacing manual feedback with automated evaluation metrics.
@@ 4.2 Peer Review
Peer review serves as a critical mechanism for improving the quality of academic manuscripts through feedback and evaluation, forming the cornerstone of quality control in scientific research. However, the process is hindered by its slow pace, high time consumption, and increasing strain due to the growing academic workload  (Lin et al., 2023a; Kousha and Thelwall, 2024; Thelwall and Yaghi, 2024). To address these challenges and enhance manuscript quality, researchers have investigated the application of AI in peer review  (Yuan et al., 2022; Liu and Shah, 2023; Niu et al., 2023; Kuznetsov et al., 2024). Peer review can be categorized into two main types: paper review generation and meta-review generation.
@@
In meta-review generation, chairs are tasked with identifying a paper’s core contributions, strengths, and weaknesses while synthesizing expert opinions on manuscript quality. Meta-reviews are conceptualized as abstractions of comments, discussions, and paper abstracts  (Li et al., 2023).  Santu et al. (2024) investigated the use of LLMs for automated meta-review generation, while  Zeng et al. (2023) proposed a guided, iterative prompting approach. MetaWriter  (Sun et al., 2024) utilized LLMs to extract key reviewer arguments, whereas GLIMPSE  (Darrin et al., 2024) and  Kumar et al. (2023) focused on reconciling conflicting statements to ensure fairness. Additionally,  Li et al. (2024b) introduced a three-layer sentiment consolidation framework for meta-review generation, and PeerArg  (Sukpanichnant et al., 2024) integrated LLMs with knowledge representation to address subjectivity and bias via a multiparty argumentation framework (MPAF).
@@ 6 Tools
To accelerate the research workflow, we have curated a collection of tools designed to support various stages of the research process, with their applicability specified for each stage. To ensure practical relevance, our selection criteria emphasize tools that are publicly accessible or demonstrate significant influence on GitHub. A comprehensive overview of these tools is presented in Table LABEL:tab:tools.
@@ 7.1 Integration of Diverse Research Tasks
Many existing studies on AI for research remain focused narrowly within their respective domains, often neglecting related technologies and potentially undermining overall outcomes. However, the research process is inherently an integrated pipeline comprising interdependent stages. Therefore, we propose that researchers strive to bridge diverse fields, either by combining technologies or harmonizing workflows. For instance, meta-review generation could be integrated with scientific claim verification, experiment verification could be linked with hypothesis formulation (Yuan et al., 2025), and research paper recommendation systems could be connected with manuscript writing processes (Gu and Hahnloser, 2023). Furthermore, some studies have begun to emphasize the development of systems capable of spanning multiple stages of the research process (Jansen et al., 2024; Weng et al., 2024; Yu et al., 2024a).
@@ 7.2 Integration with Reasoning-Oriented Language Models
Research is a process that places significant emphasis on logic and reasoning. Theorem proving serves as a subtask within logical reasoning (Li et al., 2024e), while hypothesis generation is widely recognized as the primary form of reasoning employed by scientists when observing the world and proposing hypothesis to explain these observations (Yang et al., 2024b). Experiment verification, in turn, demands a high degree of planning capability from models (Kambhampati et al., 2024). Recent advancements in reasoning-oriented language models, such as OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025), have substantially enhanced the reasoning abilities of these models. Consequently, we posit that integrating reasoning language models with reasoning tasks is a promising future direction. This prediction was validated by experiments conducted by  Schmidgall et al. (2025) using o1-Preview.
@@ 8 Conclusion
This paper provides a systematic survey of existing research on AI for research, offering a comprehensive review of the advancements in the field. Within each category, we offer detailed descriptions of the associated subfields. Furthermore, we analyze potential future research directions and address the challenges that remain unresolved. To facilitate researchers’ exploration of AI for research and enhance workflow efficiency, we provide a summary of relevant benchmarks and tools.
@@
We have made our best effort, but there may still be some limitations. On one hand, due to page limitations, we can only provide a brief summary of each method without exhaustive technical details. On the other hand, given the widespread exploration by researchers across disciplines on applying AI to their work, and our focus on articles published after 2022, it is possible that some important contributions may have been overlooked. Additionally, to prioritize areas that closely simulate the human research process, we excluded certain domains that could also be classified under AI for research.

### v2

@@ Abstract
Research is a fundamental process driving the advancement of human civilization, yet it demands substantial time and effort from researchers. In recent years, the rapid development of artificial intelligence (AI) technologies has inspired researchers to explore how AI can accelerate and enhance research. To monitor relevant advancements, this paper presents a systematic review of the progress in this domain. Specifically, we organize the relevant studies into three main categories: hypothesis formulation, hypothesis validation, and manuscript publication. Hypothesis formulation involves knowledge synthesis and hypothesis generation. Hypothesis validation includes the verification of scientific claims, theorem proving, and experiment validation. Manuscript publication encompasses manuscript writing and the peer review process. Furthermore, we identify and discuss the current challenges faced in these areas, as well as potential future directions for research. Finally, we also offer a comprehensive overview of existing benchmarks and tools across various domains that support the integration of AI into the research process. We hope this paper serves as an introduction for beginners and fosters future research.
@@
Zekun Zhou1, Xiaocheng Feng1,2thanks: Corresponding Author, Lei Huang1, Xiachong Feng3, Ziyun Song1, Ruihan Chen1, Liang Zhao1, Weitao Ma1, Yuxuan Gu1, Baoxin Wang4, Dayong Wu4, Guoping Hu4, Ting Liu1, Bing Qin1,2 1 Harbin Institute of Technology, Harbin, China  2 Peng Cheng Laboratory, Shenzhen, China 3 The University of Hong Kong, China  4 iFLYTEK Research, China {zkzhou,xcfeng,lhuang,zysong,rhchen,lzhao,wtma,yxgu,tliu,qinb}@ir.hit.edu.cn fengxc@hku.hk{bxwang2,dywu2,gphu}@iflytek.com
@@ 1 Introduction
Comparing with Existing Surveys Although Luo et al. (2025) reviews the application of AI in research, it predominantly focuses on LLMs, while neglecting the knowledge synthesis that precedes hypothesis generation and the theoretical validation of hypothesis. Other surveys concentrate on more specific areas, such as paper recommendation (Beel et al., 2016; Bai et al., 2019; Kreutz and Schenkel, 2022), scientific literature review (Altmami and Menai, 2022), hypothesis generation (Kulkarni et al., 2025), scientific claim verification (Vladika and Matthes, 2023; Dmonte et al., 2024), theorem proving (Li et al., 2024e), manuscript writing (Li and Ouyang, 2024), and peer review  (Lin et al., 2023a; Kousha and Thelwall, 2024). Additionally, certain surveys emphasize the application of AI in scientific domains (Zheng et al., 2023b; Zhang et al., 2024d; Gridach et al., 2025).
Contributions Our contributions can be summarized as follows: (1) We align the relevant fields with the research process of human researchers, systematically integrating and extending these aspects while primarily focusing on the research process itself. (2) We introduce a meticulous taxonomy (shown in Figure 2). (3) We provide a summary of tools that can assist in the research process. (4) We discuss new frontiers, outline their challenges, and shed light on future research.
@@
Formulation  (§2) [ Knowledge Synthesis  (§2.1) [ Research Paper
@@
Survey Organization We first elaborate hypothesis formulation (§2), followed by hypothesis validation (§3) and manuscript publication (§4). Additionally, we present benchmarks (§5), and tools (§6) that utilized in research. Finally, we outline challenges as well as future directions (§7). In the Appendix, we provide further discussion on open questions (§A), challenges faced in different domains (§B), discusssion about relevant ethical considerations (§C), and a comparison of capabilities among different methods (§D).
@@ 2.1.2 Systematic Literature Review
Systematic literature review (SLR) constitutes a rigorous and structured methodology for evaluating and integrating prior research on a specific topic  (Webster and Watson, 2002; Zhu et al., 2023; Bolaños et al., 2024). In contrast to single-document summaries (Elhadad et al., 2005), SLR entails synthesizing information across multiple related scientific documents (Altmami and Menai, 2022). SLR can further be divided into two stages: outline generation and full-text generation (Shao et al., 2024; Agarwal et al., 2024b; Block and Kuckertz, 2024).
@@ 2.2 Hypothesis Generation
Hypothesis quality improvement has been addressed through feedback and iteration (Rabby et al., 2025), as proposed by HypoGeniC  (Zhou et al., 2024) and MOOSE  (Yang et al., 2024b). Feedback mechanisms include direct responses to hypothesis  (Baek et al., 2024), experimental outcome evaluations  (Ma et al., 2024; Yuan et al., 2025), comparison with the existing literature (Schmidgall and Moor, 2025), and automated peer review comments  (Lu et al., 2024). FunSearch (Romera-Paredes et al., 2024) generates solutions by iteratively combining the innovative capabilities of LLM with the verification capabilities of an evaluator. Beyond iterative feedback, collaborative efforts among researchers have also been recognized, leading to multi-agent hypothesis generation approaches  (Nigam et al., 2024; Ghafarollahi and Buehler, 2024). VIRSCI  (Su et al., 2024) further optimized this process by customizing knowledge for each agent. Additionally, the Nova framework  (Hu et al., 2024a) was introduced to refine hypothesis by leveraging outputs from other research as input.
@@ 3 Hypothesis Validation
In scientific research, any proposed hypothesis must undergo rigorous validation to establish its validity. In some studies, this process is also referred to as ’falsification’ (Liu et al., 2024d; Huang et al., 2025). As illustrated in Figure  4, this section explores the application of AI in verifying scientific hypothesis through three methodological components: Scientific Claim Verification, Theorem Proving, and Experiment Validation.
@@ 3.1 Scientific Claim Verification
In the verification results generation phase, MAGIC (Kao and Yen, 2024) and SERIf (Cao et al., 2024b) proposed utilizing LLMs to synthesize evidence into more comprehensive information. FactKG  (Kim et al., 2023) and  Muharram and Purwarianti (2024) structured evidence into knowledge graphs, enabling claim attribution  (Dammu et al., 2024; Wu et al., 2023). Furthermore,  Atanasova et al. (2020); Krishna et al. (2022); Pan et al. (2023a); Eldifrawi et al. (2024); Zhang et al. (2024b) advocated for generating explanatory annotations alongside experimental outcomes during the verification process. Meanwhile,  Das et al. (2023); Altuncu et al. (2023) emphasized the critical role of domain expertise in ensuring accurate verification.
@@ 3.2 Theorem Proving
Theorem proving constitutes a subtask of logical reasoning, aimed at reinforcing the validity of the underlying theory within a hypothesis  (Pease et al., 2019; Yang et al., 2023c; Li et al., 2024e).
@@ 4.1 Manuscript Writing
Early shared tasks focused on assisting researchers in writing or analyzing linguistic features (Dale and Kilgarriff, 2010; Daudaravicius, 2015). Recent advances have led to three main directions: citation text generation, related work generation, and complete manuscript generation.
@@
A subset of research on AI in scientific writing has focused on citation text generation, which addresses the academic need for referencing prior work while mitigating model inaccuracies  (Gao et al., 2023b; Gu and Hahnloser, 2023). For instance,  Wang et al. (2022b) developed an automated citation generation system by integrating manuscript content with citation graphs. However, its reliance on rigid template-based architectures led to inflexible citation formats. This limitation motivated subsequent studies to propose incorporating citation intent as a control parameter during text generation, aiming to improve contextual relevance and rhetorical adaptability  (Yu et al., 2022; Jung et al., 2022; Koo et al., 2023; Gu and Hahnloser, 2024).
@@
In contrast to citation text generation, several studies have focused on related work generation in scholarly writing, emphasizing the production of multiple citation texts and the systematic analysis of inter-citation relationships  (Li and Ouyang, 2022, 2024). The ScholaCite framework  (Martin-Boyle et al., 2024) leveraged GPT-4 to cluster citation sources and generate draft literature review sections, although it required manually provided reference lists. By contrast, the UR3WG system  (Shi et al., 2023) adopted a retrieval-augmented architecture integrated with large language models to autonomously acquire relevant references. To improve the quality of generated related work sections,  Yu et al. (2024b) utilized GNNs to model complex relational dynamics between target manuscripts and cited literature, while  Nishimura et al. (2024) initiative advocated for explicit novelty assertions regarding referenced publications.
@@
The aforementioned investigations primarily focused on specific components of scientific writing, while a study by  Lai et al. (2024) explored the progressive generation of complete manuscripts via structured workflows. The AI-Scientist system  (Lu et al., 2024) further introduced section-wise self-reflection mechanisms to enhance compositional coherence. Several studies emphasized human-AI collaborative frameworks for improving writing efficiency  (Lin, 2024; Feng et al., 2024; Ifargan et al., 2024), whereas  Tang et al. (2024a) concentrated on enabling personalized content generation in multi-author collaborative environments. Following initial manuscript drafting, subsequent text revision processes were systematically examined  (Du et al., 2022b; Jourdan et al., 2023; Dang et al., 2025). The OREO system  (Li et al., 2022) utilized attribute classification for iterative in-situ editing, while  Du et al. (2022a); Pividori and Greene (2024) incorporated researcher feedback loops for progressive text optimization. Notably,  Kim et al. (2022); Chamoun et al. (2024); D’Arcy et al. (2024b) proposed replacing manual feedback with automated evaluation metrics.
@@ 4.2 Peer Review
Peer review serves as a critical mechanism for improving the quality of academic manuscripts through feedback and evaluation, forming the cornerstone of quality control in scientific research. However, the process is hindered by its slow pace, high time consumption, and increasing strain due to the growing academic workload  (Lin et al., 2023a; Kousha and Thelwall, 2024; Thelwall and Yaghi, 2024). To address these challenges and enhance manuscript quality, researchers have investigated the application of AI in peer review  (Yuan et al., 2022; Liu and Shah, 2023; Niu et al., 2023; Kuznetsov et al., 2024; Thakkar et al., 2025). Peer review can be categorized into two main types: paper review generation and meta-review generation.
@@
In meta-review generation, chairs are tasked with identifying a paper’s core contributions, strengths, and weaknesses while synthesizing expert opinions on manuscript quality. Meta-reviews are conceptualized as abstractions of comments, discussions, and paper abstracts  (Li et al., 2023).  Santu et al. (2024) investigated the use of LLMs for automated meta-review generation, while  Zeng et al. (2023) proposed a guided, iterative prompting approach. MetaWriter  (Sun et al., 2024) utilized LLMs to extract key reviewer arguments, whereas GLIMPSE  (Darrin et al., 2024) and  Kumar et al. (2023) focused on reconciling conflicting statements to ensure fairness. Additionally,  Li et al. (2024b) introduced a three-layer sentiment consolidation framework for meta-review generation, and PeerArg  (Sukpanichnant et al., 2024) integrated LLMs with knowledge representation to address subjectivity and bias via a multiparty argumentation framework (MPAF). DeepReview (Zhu et al., 2025) generates a comprehensive meta-review by simulating expert evaluation across multiple dimensions.
@@ 6 Tools
To accelerate the research workflow, we have curated a collection of tools designed to support various stages of the research process, with their applicability specified for each stage. To ensure practical relevance, our selection criteria emphasize tools that are publicly accessible or demonstrate significant influence on GitHub. A comprehensive overview of these tools is presented in Table 10.
@@ 7.1 Integration of Diverse Research Tasks
The research process is an integrated pipeline of interdependent stages. Paper recommendation and literature review provide an AI tool with a field overview and relevant works, ensuring that hypothesis generation is informed and of higher quality. Hypothesis validation assesses feasibility both logically and practically, with results feeding back to refine the hypothesis. In manuscript writing, validated hypotheses and prior outputs serve as key inputs. Peer review evaluates the manuscript and offers feedback across modules, enabling the hypothesis generator to adjust content accordingly (Lu et al., 2024). In addition, combinations can also be made between some small fields, for instance, meta-review generation could be integrated with scientific claim verification, experiment verification could be linked with hypothesis formulation (Jansen et al., 2025; Yuan et al., 2025; Liu et al., 2024d), and research paper recommendation systems could be connected with manuscript writing processes (Gu and Hahnloser, 2023). Furthermore, some studies have begun to emphasize the development of systems capable of covering multiple stages of the research process (Jansen et al., 2024; Weng et al., 2024; Yu et al., 2024a).
@@ 7.2 Integration with Reasoning-Oriented Language Models
Research is a process that places a significant emphasis on logic and reasoning. Theorem proving serves as a subtask within logical reasoning (Li et al., 2024e), while hypothesis generation is widely recognized as the primary form of reasoning employed by scientists when observing the world and proposing hypothesis to explain these observations (Yang et al., 2024b). Experiment verification, in turn, demands a high degree of planning capability from models (Kambhampati et al., 2024). Recent advances in reasoning-oriented language models, such as OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025), have substantially enhanced the reasoning abilities of these models. Consequently, we posit that integrating reasoning language models with reasoning tasks is a promising future direction. This prediction was validated by experiments conducted by  Schmidgall et al. (2025) using o1-Preview.
@@ 8 Conclusion
This paper provides a systematic survey of existing research on AI for research, offering a comprehensive review of the advancements in the field. Within each category, we offer detailed descriptions of the associated subfields. In addition, we examine current challenges, ethical considerations, and potential directions for future research. To support researchers in exploring AI-driven research applications and enhancing workflow efficiency, we also summarize existing benchmarks and tools, accompanied by a comparative analysis of representative methods and their capabilities.
@@
We have made our best effort, but there may still be some limitations. Due to space constraints, we provide only concise summaries of each method without detailed technical elaboration. Given the rapid progress in AI and the expanding research landscape, we primarily focus on works published after 2022, with earlier studies receiving less attention. To emphasize areas that closely mimic the human research process, some topics are excluded from the main text but briefly discussed in Appendix §A. Moreover, as AI for Research is still an emerging field, the lack of standardized benchmarks and evaluation metrics hinders direct comparison. Nonetheless, we offer a comparative analysis of representative methods across domains using attribute graphs in Appendix §D.